OpenStudy (anonymous):

Evaluate: 2(6 − 3) − 1 • 6

OpenStudy (turingtest):

it's best to just type your question in the post ;)

OpenStudy (anonymous):

Evaluate: 2(6 − 3) − 1 • 6

OpenStudy (anonymous):

2(6-3)-1*6 =2(3)-6 =6-6 =0

OpenStudy (anonymous):

thankyouu